OpenStudy (anonymous):

AB ← → and BC ← → form a right angle at point B. If A = (-3, -1) and B = (4, 4), what is the equation of BC ← → ?

OpenStudy (anonymous):

x + 3y = 16 2x + y = 12 -7x − 5y = -48 7x − 5y = 48

OpenStudy (gorv):



OpenStudy (gorv):

BC is perpendicular to AB

OpenStudy (gorv):

slope of BC = -1/(slope of AB)

OpenStudy (gorv):

slope of AB=(4-(-1))/(4-(-3))

OpenStudy (gorv):

slope of AB=(4+1)/(4+3)

OpenStudy (gorv):

cal culate this and find slope of BC

OpenStudy (gorv):

and BC pass through (4,4) ....u have one point and slope ..u can do further